OpenStudy (anonymous):

Which sufficient condition have to satisfied function f:D->R and areas D and D' that...

OpenStudy (anonymous):

\[\int\limits_{D}^{}\int\limits_{}^{}f*dP=2\int\limits_{D'}^{}\int\limits_{}^{}f*dP\]